# Unleashing the Potential of the Diffusion Model in Few-shot Semantic Segmentation

### v1

@@ Abstract
The Diffusion Model has not only garnered noteworthy achievements in the realm of image generation but has also demonstrated its potential as an effective pretraining method utilizing unlabeled data. Drawing from the extensive potential unveiled by the Diffusion Model in both semantic correspondence and open vocabulary segmentation, our work initiates an investigation into employing the Latent Diffusion Model for Few-shot Semantic Segmentation. Recently, inspired by the in-context learning ability of large language models, Few-shot Semantic Segmentation has evolved into In-context Segmentation tasks, morphing into a crucial element in assessing generalist segmentation models. In this context, we concentrate on Few-shot Semantic Segmentation, establishing a solid foundation for the future development of a Diffusion-based generalist model for segmentation. Our initial focus lies in understanding how to facilitate interaction between the query image and the support image, resulting in the proposal of a KV fusion method within the self-attention framework. Subsequently, we delve deeper into optimizing the infusion of information from the support mask and simultaneously re-evaluating how to provide reasonable supervision from the query mask. Based on our analysis, we establish a simple and effective framework named DiffewS, maximally retaining the original Latent Diffusion Model’s generative framework and effectively utilizing the pre-training prior. Experimental results demonstrate that our method significantly outperforms the previous SOTA models in multiple settings.
@@ 1 Introduction
Let’s reconsider the most fundamental question in using generative models for visual perception: how to design a fine-tuning framework that can guarantee both generalization ability and precise prediction of details? Unfortunately, existing methods do not sufficiently address this challenge. The demands of the FSS task for open-set generalization and high-quality segmentation results precisely align with this challenge. Thus, our first motivation is to further address the fundamental question posed above by exploring the Diffusion Model on the FSS task.
@@ 2 Related Work
Diffusion models have shown impressive performance on visual generation tasks such as text-based image generation [1, 2], image editing [3, 4], and video generation [5, 6, 7]. Recent work explores visual priors of diffusion models for perceptual tasks. ODISE [16] uses frozen diffusion models for panoptic segmentation of any category in the wild. DVP [12], DatasetDM [17] adapt LDM to multiple tasks such as depth estimation, and semantic segmentation. Marigold [13] fine-tunes diffusion models on synthetic data for affine-invariant monocular depth estimation and achieves impressive performance. Different from the above methods, we focus on using diffusion models to model the visual correlations of multiple reference images and a target image for few-shot segmentation.
Few-shot semantic segmentation [31, 32] aims to segment target objects in an input image given a few annotated support images. Traditional FSS methods either explore prototype learning [33, 34, 35] of support images to predict query images’ masks or use pixel-level information [36, 37, 18] to exploit the support information. For example, some works [29, 30] demonstrate powerful generalization ability by unifying various segmentation tasks in an in-context learning framework. SegGPT [30] can exactly segment any semantic conception by using one or a few support images, which motivates us to explore the potential of the diffusion model for the FSS task under the in-context setting [30].
@@ 3 Preliminary
Figure 1: Overview of the DiffewS framework. (a)(b) display that query image $\mathbf{I}_{q}$, query mask $\mathbf{M}_{q}$, support image $\mathbf{I}_{s}$ and support mask $\mathbf{M}_{s}$ are all encoded by VAE into latent variables $\mathbf{z}_{q}$, $\mathbf{z}_{mq}$, $\mathbf{z}_{s}$, $\mathbf{z}_{ms}$, respectively, where $\mathbf{z}_{q}$ and $\mathbf{z}_{mq}$ are concatenated to input into UNet. (c) demonstrates the DiffewS fintuning protocol, where the query image branch performs the original Self-attention while the support Image branch facilitates interaction between query image and support image by executing KV Fusion Self-Attention (FSA). (d) elucidates the detailed implementation of FSA, acquiring information from support images by concatenating the query and key features.
@@ 4.2 Injection of support mask information
    Attention Mask $\mathbf{M}_{s}$ can serve as an attention mask to control self-attention, so that only $\mathbf{K}_{s}$ in the masked region can be accessed by $\mathbf{Q}_{q}$. Since the feature map sizes of different layers are different, we need to resize $\mathbf{M}_{s}$ to fit the dimensions of each layer.
@@ 4.4 Exploration of generation process
We explore the mask generation pipeline starting from MN2M. As shown in Figure 4(c), MN2N achieves 15.2% mIoU. Then, we change MN2M into MI2M keeping same variance $\beta_{1}=(0.00085,0.012)$. The performance has improved by 4.7% mIoU. Both methods achieve bad performance. We hypothesize that this is because adding a very small noise or image to the binary mask during the training process and then predicting it does not lead to a challenging task compared with diffusion pre-training. The binary mask is simpler than natural images, the mask latent after adding noise can still easily distinguish the foreground and the background. This results in serious information leakage in the training process of UNet, leading to poor performance. To verify this hypothesis, we increase the variance of MI2M from $\beta_{1}=(0.00085,0.012)$ to $\beta_{2}=(0.0272,0.384)$. The performance has significantly improved by 23.3% mIoU. To fully increase the difficulty of training, we convert MI2M into OI2M, which does not introduce any ground-truth information into the input of the UNet during training. In addition, OI2M reduces the number of iterations to 1, which significantly improves the predictive efficiency of the network. Figure 4(c) shows OI2M achieves best performance. Therefore, we choose OI2M as the mask generation pipeline.
@@ 5 Experiment
Implementation details We initialize our model with Stable Diffusion 2.1 [2]. The Adam optimizer is used with a weight decay set at 0.01 and a learning rate of 1e-5, coupled with a linear schedule. In terms of data augmentation, our methodology only involves resizing the input image directly to 512x512. No additional data augmentation occurs. Under the strict few-shot setting, the model undergoes training on four V100 GPUs. With the gradient accumulation set at 4, the total batch size comes to 16. Training carries out for 10,000 iterations, typically requiring six hours. For in-context setting, since the training set is larger, we keep other hyperparameters consistent with the strict few-shot setting, and adjust the total training iterations to 30000 iterations. Lastly, our ablation experiments are validated on Fold0 of COCO-20i [43]. The training took place on a single 4090 GPU, with a gradient accumulation set at 4, which brought the total batch size to 4. The training, which consisted of 10,000 iterations, took roughly 11 hours.
@@
Table 2: Results of strict few-shot semantic segmentation on COCO-20i.
@@ 5.1 In-context setting
We first compare DiffewS with other generalist models such as Painter [29], SegGPT [29], PerSAM-F[49], and Matcher [22] as well as specialist models like HSNet [18], VAT [47], FPTrans [48]. Regarding the specialist models, we directly refer to the results presented within the SegGPT [30] and Matcher [22] research papers. These specialist models are also trained on the test categories from COCO [44] and PASCAL VOC [46]. We employ COCO-20i [43] and PASCAL-5i [31] to validate the in-domain performance of DiffewS. Remarkably, on COCO, DiffewS achieves a 1-shot score of 71.3, considerably exceeding the generalist model SegGPT (+15.2) and specialist model FPTrans (+14.8), both trained with in-domain data. DiffewS furthermore significantly outperforms SAM-based models PerSAM-F (+47.8) and Matcher (+18.6). On PASCAL-5i, DiffewS records 88.3 in 1-shot, clearly surpassing SegGPT (+5.1) and Matcher (+20.4). These results evidence that DiffewS effectively utilizes the prior of Stable Diffusion, unlocking the full potential of Stable Diffusion in segmentation. Furthermore, out-of-domain examination on LVIS-92i [22] underpins the generalization ability of DiffewS. In this setting, DiffewS registers 31.4 in 1-shot and 35.4 in 5-shot, markedly outperforming other generalist models, aside from Matcher. It is worth mentioning that Matcher simultaneously utilizes two Foundation models (SAM [21] and DINO V2 [23] ), and SAM itself is pre-trained on an exhaustive, finely annotated segmentation dataset. On the other hand, DiffewS undergoes fine-tuning on a relatively smaller quantity of segmentation data for limited iterations, still delivering performance that rivals Matcher. This indicates that using the paradigm of DiffewS, there is potential to achieve significant breakthroughs in the segmentation field if further trained on larger-scale segmentation data. It should be noted that the improvement of DiffewS in 5-shot is not significant, with a 4.0 distinct improvement only on LVIS-92i. This might be due to the presence of many small objects in the support images of LVIS, so increasing the number of support images can alleviate this problem. Conversely, the DiffewS 5-shot performance on PASCAL-5i [31] is slightly deficient compared to the 1-shot. This could be ascribed to the presence of relatively larger and more simplistic objects within PASCAL VOC’s support images, inputting more images might interfere with the original architecture of the model. In the Section A.6, details of how to extend the inference to a 5-shot scenario are provided, and a discussion about this phenomenon will occur in the next section Section 5.2.
@@ 5.2 Strict few-shot setting
We also undertake validation of DiffewS under the standard few-shot setting, comparing it with other specialist models such as HSNet [18], CyCTR [50], VAT [47], BAM [51], HDMNet [20], and DCAMA [19]. For the one-shot setting, the average performance of DiffewS across all four folds attains 51.2, surpassing the current state-of-the-art (SOTA) model DCAMA, scoring 50.9 mIoU. Worth mentioning is that DCAMA relies on a highly complex additional block, whereas DiffewS entirely utilizes the generative framework of UNet. In terms of the efficiency of convergence, DiffewS necessitates just a 30000-iteration training, in contrast to both DCAMA and HSNet which require training spanning hundreds of epochs, typically costing several days. This demonstrates the successful employment of Stable Diffusion priors by DiffewS, thereby securing impressive performance without requiring extended periods of fine-tuning. In the five-shot setting, the average performance across four folds reaches 56.0, higher than all other models aside from DCAMA. Currently, DiffewS primarily focuses on the 1-shot situation lacking specific optimizations for the 5-shot scenario in its training and inference systems. This explains why DiffewS is at present marginally inferior to DCAMA.
@@ 5.3 Visualization
As shown in Figure 5, DiffewS effectively segments categories not in the training set, such as slippers and aprons. It also accurately segments objects of different styles and smaller items, demonstrating strong generalization capabilities. In some cases, DiffewS even achieves more accurate results than GT.
@@ A.4 Post processing
Our experiments (see Table 3) have shown that the relative threshold method achieved better results on COCO-20i [43] fold0. The optimal $\tau_{r}$ is 0.25.
@@ A.5 More ablation studies
As shown in 4, the Multiplication method directly applied to RGB images achieved better results. However, the overall disparity is not significant.
@@
A.6 1-shot to n-shot
A.7 Other visualization

### v3

@@ Abstract
The Diffusion Model has not only garnered noteworthy achievements in the realm of image generation but has also demonstrated its potential as an effective pretraining method utilizing unlabeled data. Drawing from the extensive potential unveiled by the Diffusion Model in both semantic correspondence and open vocabulary segmentation, our work initiates an investigation into employing the Latent Diffusion Model for Few-shot Semantic Segmentation. Recently, inspired by the in-context learning ability of large language models, Few-shot Semantic Segmentation has evolved into In-context Segmentation tasks, morphing into a crucial element in assessing generalist segmentation models. In this context, we concentrate on Few-shot Semantic Segmentation, establishing a solid foundation for the future development of a Diffusion-based generalist model for segmentation. Our initial focus lies in understanding how to facilitate interaction between the query image and the support image, resulting in the proposal of a KV fusion method within the self-attention framework. Subsequently, we delve deeper into optimizing the infusion of information from the support mask and simultaneously re-evaluating how to provide reasonable supervision from the query mask. Based on our analysis, we establish a simple and effective framework named DiffewS, maximally retaining the original Latent Diffusion Model’s generative framework and effectively utilizing the pre-training prior. Experimental results demonstrate that our method significantly outperforms the previous SOTA models in multiple settings. Code is released at: https://github.com/aim-uofa/DiffewS
@@ 1 Introduction
Let us reconsider the most fundamental question in using generative models for visual perception: how to design a fine-tuning framework that can guarantee both generalization ability and precise prediction of details? Unfortunately, existing methods do not sufficiently address this challenge. The demands of the FSS task for open-set generalization and high-quality segmentation results precisely align with this challenge. Thus, our first motivation is to further address the fundamental question posed above by exploring the Diffusion Model on the FSS task.
@@ 2 Related Work
Diffusion models have shown impressive performance on visual generation tasks such as text-based image generation [1, 2], image editing [3, 4], and video generation [5, 6, 7]. Current research on leveraging Diffusion models to enhance visual perception tasks mainly focuses on two directions: one is the direct use of diffusion models to generate images, aiming to address the issue of insufficient data, such as instance segmentation [31, 32, 33], semantic segmentation [34], few-shot segmentation [35] and so on. Another direction is to transfer features from Diffusion models to other visual tasks, which aligns with the research direction of this paper.
ODISE [16] uses frozen diffusion models for panoptic segmentation of any category in the wild. DVP [12], DatasetDM [17], GenPercept [36], Geowizard [37] adapt LDM to multiple tasks such as depth estimation, semantic segmentation, and surface normal. Marigold [13] fine-tunes diffusion models on synthetic data for affine-invariant monocular depth estimation and achieves impressive performance. Different from the above methods, we focus on using diffusion models to model the visual correlations of multiple reference images and a target image for few-shot segmentation. The most related work to this paper is a concurrent study [38], which focuses on utilizing diffusion models for in-context segmentation. However, it disrupts the original U-Net structure and the priors of the diffusion model to some extent. In contrast, our work offers a more comprehensive and systematic analysis of applying diffusion models to Few-Shot Semantic Segmentation tasks.
Few-shot semantic segmentation [39, 40] aims to segment target objects in an input image given a few annotated support images. Traditional FSS methods either explore prototype learning [41, 42, 43] of support images to predict query images’ masks or use pixel-level information [44, 45, 18] to exploit the support information. For example, some works [29, 30, 46] demonstrate powerful generalization ability by unifying various segmentation tasks in an in-context learning framework. SegGPT [30] can exactly segment any semantic conception by using one or a few support images, which motivates us to explore the potential of the diffusion model for the FSS task under the in-context setting [30].
@@ 3 Preliminary
Figure 1: Overview of the DiffewS framework. (a)(b) display that query image $\mathbf{I}_{q}$, query mask $\mathbf{M}_{q}$, support image $\mathbf{I}_{s}$ and support mask $\mathbf{M}_{s}$ are all encoded by VAE into latent variables $\mathbf{z}_{q}$, $\mathbf{z}_{mq}$, $\mathbf{z}_{s}$, $\mathbf{z}_{ms}$, respectively, where $\mathbf{z}_{q}$ and $\mathbf{z}_{mq}$ are concatenated to input into UNet. (c) demonstrates the DiffewS fintuning protocol (d) elucidates the detailed implementation of FSA, acquiring information from support images by concatenating the query and key features.
@@ 4.2 Injection of support mask information
    Attention Mask $\mathbf{M}_{s}$ can serve as an attention mask to control self-attention so that only $\mathbf{K}_{s}$ in the masked region can be accessed by $\mathbf{Q}_{q}$. Since the feature map sizes of different layers are different, we need to resize $\mathbf{M}_{s}$ to fit the dimensions of each layer.
@@ 4.4 Exploration of generation process
We explore the mask generation pipeline starting from MN2M. As shown in Figure 4(c), MN2N achieves 15.2% mIoU. Then, we change MN2M into MI2M keeping same variance $\beta^{1}=(0.00085,0.012)$, respectively representing the initial and final values of $\beta$ in the DDIM scheduler. The performance has improved by 4.7% mIoU. However, despite the improvement, both methods exhibit suboptimal performance. We hypothesize that this is because adding a very small noise or image to the binary mask during the training process and then predicting it does not lead to a challenging task compared with diffusion pre-training.
We hypothesize that the suboptimal performance is due to the minimal noise or image added to the binary mask during training, which results in an insufficiently challenging task compared to diffusion pre-training. The binary mask is inherently simpler than natural images, and even after adding noise, the latent mask can still easily distinguish between the foreground and background. This simplicity causes significant information leakage during UNet training, ultimately leading to poor performance.
To verify this hypothesis, we increase the variance of MI2M from $\beta^{1}=(0.00085,0.012)$ to $\beta^{2}=(0.0272,0.384)$. The performance has significantly improved by 23.3% mIoU. To fully increase the challenge of training, we convert MI2M into OI2M, which does not introduce any ground-truth information into the input of the UNet during training. Additionally, OI2M reduces the number of iterations to one, significantly boosting the network’s predictive efficiency. As shown in Figure 4(c), OI2M achieves the best performance, making it the preferred choice for the mask generation pipeline.
4.5 1-shot to N-shot
So far, we have primarily explored the training and inference processes specifically designed for 1-shot scenarios. A natural question arises: can this framework be extended to n-shot settings? To address this, we first present the simplest and most straightforward method for adaptation, which requires only minor modifications during the inference phase to accommodate n-shot tasks.
While the aforementioned solutions enable N-shot inference, their performance does not match that of state-of-the-art (SOTA) models. This discrepancy primarily arises because the model receives only a single support image during the training phase, which leads to inconsistencies when transitioning to the inference phase with 5-shot or 10-shot configurations.
To address this issue, we explore improvements from both the inference and training perspectives. From the perspective of inference, transitioning from 1-shot to N-shot involves concatenating the keys and values of additional support samples, which significantly increases the number of keys and values processed during inference. To address this, we implement random sampling of the keys and values from the support samples during inference, ensuring that their quantity matches that of the training phase (see Table 6). Another more straightforward idea is to introduce multiple support samples during the training phase. In this way, the model can learn how to utilize multiple support images during training. we randomly select 1 to N support samples as input using KV Fusion in Equation 8 during a single training iteration (see Table 7).
Our experiments demonstrate that improvements during the training phase are more effective than those during the inference phase. Therefore, we include the results of the model with training phase improvements in Table 2.
@@ 5 Experiment
Implementation details We initialize our model with Stable Diffusion 2.1 [2]. The Adam optimizer is used with a weight decay set at 0.01 and a learning rate of 1e-5, coupled with a linear schedule. In terms of data augmentation, our methodology only involves resizing the input image directly to 512x512. No additional data augmentation occurs. Under the strict few-shot setting, the model undergoes training on four V100 GPUs. With the gradient accumulation set at 4, the total batch size comes to 16. Training carries out for 10,000 iterations, typically requiring six hours. For in-context setting, since the training set is larger, we keep other hyperparameters consistent with the strict few-shot setting, and adjust the total training iterations to 30000 iterations. Lastly, our ablation experiments are validated on Fold-0 of COCO-20i [52]. The training took place on a single 4090 GPU, with a gradient accumulation set at 4, which brought the total batch size to 4. The training, which consisted of 10,000 iterations, took roughly 11 hours.
@@
Table 2: Results of strict few-shot semantic segmentation on COCO-20i. DiffewS-n represents using training time improvements for N-shot.
@@ 5.1 In-context setting
We first compare DiffewS with other generalist models such as Painter [29], SegGPT [29], PerSAM-F[58], and Matcher [22] as well as specialist models like HSNet [18], VAT [56], FPTrans [57]. Regarding the specialist models, we directly refer to the results presented within the SegGPT [30] and Matcher [22] research papers. These specialist models are also trained on the test categories from COCO [53] and PASCAL VOC [55]. We employ COCO-20i [52] and PASCAL-5i [39] to validate the in-domain performance of DiffewS. Remarkably, on COCO, DiffewS achieves a 1-shot score of 71.3, considerably exceeding the generalist model SegGPT (+15.2) and specialist model FPTrans (+14.8), both trained with in-domain data. DiffewS furthermore significantly outperforms SAM-based models PerSAM-F (+47.8) and Matcher (+18.6). On PASCAL-5i, DiffewS records 88.3 in 1-shot, clearly surpassing SegGPT (+5.1) and Matcher (+20.4). These results evidence that DiffewS effectively utilizes the prior of Stable Diffusion, unlocking the full potential of Stable Diffusion in segmentation. Furthermore, out-of-domain examination on LVIS-92i [22] underpins the generalization ability of DiffewS. In this setting, DiffewS registers 31.4 in 1-shot and 35.4 in 5-shot, markedly outperforming other generalist models, aside from Matcher. It is worth mentioning that Matcher simultaneously utilizes two Foundation models (SAM [21] and DINO V2 [23] ), and SAM itself is pre-trained on an exhaustive, finely annotated segmentation dataset. On the other hand, DiffewS undergoes fine-tuning on a relatively smaller quantity of segmentation data for limited iterations, still delivering performance that rivals Matcher. This indicates that using the paradigm of DiffewS, there is potential to achieve significant breakthroughs in the segmentation field if further trained on larger-scale segmentation data. It should be noted that the improvement of DiffewS in 5-shot is not significant, with a 4.0 distinct improvement only on LVIS-92i. This might be due to the presence of many small objects in the support images of LVIS, so increasing the number of support images can alleviate this problem. Conversely, the DiffewS 5-shot performance on PASCAL-5i [39] is slightly deficient compared to the 1-shot. This could be ascribed to the presence of relatively larger and more simplistic objects within PASCAL VOC’s support images, inputting more images might interfere with the original architecture of the model. In this case, we do not apply the improvement strategies discussed in Section 4.5, therefore, the relatively weaker performance in the 5-shot scenario is reasonable.
@@ 5.2 Strict few-shot setting
We also undertake validation of DiffewS under the standard few-shot setting, comparing it with other specialist models such as HSNet [18], CyCTR [59], VAT [56], BAM [60], HDMNet [20], and DCAMA [19]. For the one-shot setting, the average performance of DiffewS across all four folds attains 51.2, surpassing the current state-of-the-art (SOTA) model DCAMA, scoring 50.9 mIoU. Worth mentioning is that DCAMA relies on a highly complex additional block, whereas DiffewS entirely utilizes the generative framework of UNet. In terms of the efficiency of convergence, DiffewS necessitates just a 30000-iteration training, in contrast to both DCAMA and HSNet which require training spanning hundreds of epochs, typically costing several days. This demonstrates the successful employment of Stable Diffusion priors by DiffewS, thereby securing impressive performance without requiring extended periods of fine-tuning. In the five-shot setting, the average performance across four folds reaches 56.0, higher than all other models aside from DCAMA. Currently, DiffewS primarily focuses on the 1-shot situation lacking specific optimizations for the 5-shot scenario in its training and inference systems. This explains why DiffewS is at present marginally inferior to DCAMA. Furthermore, when employing our proposed training improvement strategy, DiffewS-n outperforms other models in both the 1-shot and 5-shot settings.
@@ 5.3 Visualization
As shown in Figure 5, DiffewS effectively segments categories not in the training set, such as slippers and aprons. It also accurately segments objects of different styles and smaller items, demonstrating strong generalization capabilities. In some cases, DiffewS even achieves more accurate results than ground-truth human annotation.
In addition, DiffewS demonstrates impressive results in various cross-style segmentation tasks and small object segmentation cases (see Figure 6). We hypothesize that DiffewS’s exceptional generalization ability stems from its extensive utilization of prior knowledge from diffusion models. However, DiffewS also struggles with certain challenging cases, we also present several failure cases in Figure 7 and categorize the reasons for these failures.
@@ A.1 Discussion
On the other hand, we believe that our method is not limited to FSS. Our framework has the potential to unify few-shot segmentation and open vocabulary segmentation by leveraging prompts from different modalities, as some work [61, 62, 63] has already proven the possibility.
@@ A.4 Post processing
Our experiments (see Table 3) have shown that the relative threshold method achieved better results on COCO-20i [52] fold-0. The optimal $\tau_{r}$ is 0.25.
@@ A.5 More ablation studies
As shown in Table 4, the Multiplication method directly applied to RGB images achieved better results. However, the overall disparity is not significant.
@@
A.6 Other visualization